OpenStudy (anonymous):

the surface areas of two similar solids are 311ft^2 and 1,307ft^2. the volume of the larger solid is 1,755ft^3. what is the volume of the smaller solid?

OpenStudy (anonymous):

1,307/311=1755/x 1307x=1755(311) 1307x=545805 x=417.6013770

OpenStudy (anonymous):

417.6ft^3

OpenStudy (anonymous):

A) 961ft^3 B) 1755ft^3 C) 789ft^3 D) 288ft^3

OpenStudy (anonymous):

these are my answer choices